# Clustering and Runaway Merging in a Primordial Black Hole Dominated Universe

### v1

@@ I Introduction
First postulated by Hawking in 1971 [1], primordial black holes (PBH) may have formed in the early universe from large density fluctuations on small scales; see Ref. [2] for a review. If the PBH mass scale falls within $10^{17}\,\mathrm{g}<m<10^{23}\,\mathrm{g}$, PBHs can viably accommodate the total dark matter abundance in our universe [3, 4]; for $m>10^{23}\,\mathrm{g}$, these objects are subject to various astrophysical bounds and can account for a small fraction of dark matter. The cosmological abundance of lighter PBHs with mass $<10^{17}\,\mathrm{g}$ is generically subject to stringent cosmological limits from the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) [5].
@@
Earlier work has considered clusters of cosmologically stable PBH in the later universe after matter-radiation equality [20, 21, 22, 23], or focused on PBH mergers in the early Universe without cluster formation (the “quasi-linear regime”) [10, 21, 24, 25, 26, 27, 28, 7, 8, 9], while Refs. [28, 29] explored the formation of heavier PBH from strongly clustered lighter ones (clusteringenesis), but did not specify a mechanism for this enhanced initial clustering at PBH formation. Ref. [30] studied PBH formation from the growth of density perturbations in PBH domination but did not consider mergers in PBH clusters. Black hole mergers in clusters have also been studied in the context of building supermassive black holes in galaxies [31, 32, 33]. Our analysis is complementary as we consider evaporating PBHs with $m<10^{9}\,\mathrm{g}$ and calculate their clustering behavior in the early universe in a BHD era, which has not been studied before.
@@ III.1 PBH Cluster Formation
$\mu(M)=\begin{cases}\displaystyle 1-\frac{4}{7}\frac{M}{M_{H,i}}&M<M_{H,i}\vspace{0.3em}\\ \displaystyle\frac{3}{7}{\left(\frac{M}{M_{H,i}}\right)}^{-4/3}&M>M_{H,i}\end{cases}$ (21)
@@ III.2 PBH Binary Capture
where $t_{\mathrm{cap,ins}}$ are respectively the capture and inspiral timescales and the 2b/3b labels refer to two and three body processes. Thus, on average, 2-body interactions lead to a quicker capture, inspiral, and merger than 3-body interactions, so for the remainder of this work we can neglect 3-body interactions, which will not contribute significantly to the PBH merger rate during BHD. In the remainder of this paper, we focus on the 2-body interaction channel.
@@ IV.3 Combined Treatment
[…]
@@ IV.4 Merged Relics
$N_{\mathrm{col}}^{\mathrm{max}}\approx 10^{10}\left\{\begin{array}[]{l}\displaystyle 12\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!12/7}{\left(\frac{t_{i}}{10^{-20}\,\mathrm{s}}\right)}^{\!-4/7}\vspace{0.3em}\\ \displaystyle 6.5\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!8/11}\end{array}\right.\!\!\!,$ (55)
@@
$m_{\mathrm{relic}}^{\mathrm{max}}\approx 10^{15}\,\mathrm{g}\left\{\begin{array}[]{l}\displaystyle 24\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!3.53}{\left(\frac{t_{i}}{10^{-20}\,\mathrm{s}}\right)}^{\!-1.27}\vspace{0.3em}\\ \displaystyle 1.4\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!1.35}\end{array}\right.\!\!\!,$ (56)
@@
Figure 2: A summary of important timescales in our scenario for a range of initial cluster sizes $N_{i}$, for two choices of free parameters $m_{i},t_{i}$. Relevant cosmic events are labeled at the top, including BHD, BBN, CMB, and the present time. The purple
@@
line gives the cluster size corresponding to the mass contained within the horizon. The green
@@
curve shows the size of the PBH cluster which can form at a given cosmic time. At a later time, indicated by the dark blue
@@
runaway merger curve, PBH clusters of a given size collapse and merge to a single BH. Finally, this merged relic evaporates at the time of the blue
@@
Figure 3: Average PBH mass as a function of initial cluster PBH number $N_{i}$ at different times, for two choices of initial parameters $m_{i}$ and $t_{i}$. Initially at cluster formation ($t=t_{\mathrm{cl}}$), the average mass is $m_{i}$ (green
@@
line), but eventually some clusters undergo runaway mergers and their average mass grows, finally reaching the mass shown by the dark blue
@@
curve shows the current remaining relic masses after the cosmologically unstable ones have evaporated away by the present day ($t=t_{0}$). Note that for the parameters on the right, all relics evaporate before the present day. The top panel on either side shows the fractional distribution of PBHs in each cluster size at the time of PBH evaporation $\tau$ in gray
@@
[…]
@@ V Observables and Constraints
Figure 4: The mass function of merged relic PBHs for two choices of free parameters. Initially, all the PBHs have the same mass $m_{i}$, shown in green
$\blacksquare$
. The blue
$\blacksquare$
and red
$\blacksquare$
curves show the initial and present-day mass functions of merged relics, which clearly shows the mass growth from runaway mergers. Superimposed in blue
$\blacksquare$
@@ V.1 Merged Relics as Dark Matter Components
The BHD dynamics outlined in Sec. III predict that for many choices of initial parameters ($m_{i},t_{i}$), there is an appreciable population of merged relics with significantly larger masses than the original PBH population that drives BHD, $m_{\mathrm{relic}}\gg m_{i}$. Consequently, these objects are longer lived and generically survive past BBN to affect the later universe. There are two qualitatively distinct scenarios for relic masses:
  • Evaporating Relics: For intermediate masses $10^{9}\,\mathrm{g}<m_{\mathrm{relic}}<10^{15}\,\mathrm{g}$, the relics evaporate after BHD and are not constrained by late-time observations. However, because of the energy they inject while evaporating, they do face stringent limits from the successful predictions of standard BBN and CMB cosmology [4, 3].
  • Metastable Relics: For $m_{\mathrm{relic}}>10^{15}\,\mathrm{g}$, relics can be cosmologically metastable and account for an appreciable fraction of the dark matter density. This population is constrained by late-time observations including lensing, gravitational waves, and particle fluxes from Hawking evaporation, although there is an “asteroid mass” window between $10^{17}-10^{23}\,\mathrm{g}$ with no current constraints [4, 3].
@@ V.2 Gravitational Waves
where $g_{*,S}$ is the number of entropic degrees of freedom, an RH label denotes a quantity at reheating ($t=\tau$), a 00 label represents a present day quantity, and $m$ is the characteristic mass of the PBHs that undergo runaway mergers. Although the peak frequency in Eq. (61) is far outside of any known observational technique, as noted in Ref. [10], the infrared tail of this stochastic background might be within the reach of the Big Bang Observer [82, 83].
@@ VI Conclusion
However, if the initial PBHs have a shot noise power spectrum, we find that relics cannot accommodate the full dark matter abundance. We find that, by assembling relic PBHs in the viable dark matter window ($m_{\mathrm{relic}}\sim 10^{17}-10^{23}\,\mathrm{g}$) with the observed DM abundance, there is an irreducible tail of evaporating relics at lower masses ($10^{9}-10^{15}\,\mathrm{g}$) whose abundance is excluded by the BBN and CMB limits. Thus, under our minimal assumptions, this mechanism cannot account for any viable present day PBH abundance. It remains to be seen whether a different ansatz for the initial PBH power spectrum (e.g. more initial clustering than shot noise predicts) could allow for a viable dark matter formation scenario and we leave such analyses for future work.
Finally, we note that our results rely on several assumptions about PBH cluster evolution discussed in Sec. IV.5. Notably, we assume that gravitational bremsstrahlung, cluster-cluster interactions, and relativistic effects (among others) do not significantly affect our results. While we have presented plausible arguments for why our results should not be affected in most of these cases, many of these effects are not analytically tractable and may, nonetheless, allow for different conclusions to hold; especially when multiple effects are combined. Thus, in order to robustly characterize cluster evolution and relic formation during BHD, a full $N$-body simulation will ultimately be necessary. Running such a simulation is beyond the scope of the present work, but we strongly encourage future efforts to better understand this simple, yet physically rich cosmology.
@@ Appendix A Shot Noise Power Spectrum
$\begin{split}\delta_{\mathrm{BH}}(\mathbf{k})&=\int d^{3}\mathbf{x}\,\delta_{\mathrm{BH}}(\mathbf{x})e^{-i\mathbf{k}\cdot\mathbf{x}}\\ &=\int d^{3}\mathbf{x}\,\left(\frac{1}{\bar{n}_{\mathrm{BH}}}\sum_{i}\delta_{D}^{(3)}(\mathbf{x}-\mathbf{x}_{i})-1\right)e^{-i\mathbf{k}\cdot\mathbf{x}}\\ &=\frac{1}{\bar{n}_{\mathrm{BH}}}\sum_{i}\int d^{3}\mathbf{x}\,\delta_{D}^{(3)}(\mathbf{x}-\mathbf{x}_{i})e^{-i\mathbf{k}\cdot\mathbf{x}}-\int d^{3}\mathbf{x}\,e^{-i\mathbf{k}\cdot\mathbf{x}}\\ &=\frac{1}{\bar{n}_{\mathrm{BH}}}\sum_{i}e^{-i\mathbf{k}\cdot\mathbf{x}_{i}}-(2\pi)^{3}\delta_{D}^{(3)}(\mathbf{k})\\ \end{split}$ (64)

### v2

@@ I Introduction
First postulated by Hawking in 1971 [1], primordial black holes (PBH) may have formed in the early universe from large density fluctuations on small scales; see Ref. [2] for a review. If the PBH mass scale falls within $10^{17}\,\mathrm{g}<m<10^{23}\,\mathrm{g}$, PBHs can viably accommodate the total dark matter abundance in our universe [3, 4]; for $m>10^{23}\,\mathrm{g}$, these objects are subject to various astrophysical bounds and can account for a small fraction of dark matter. The cosmological abundance of lighter PBHs with mass $10^{17}\,\mathrm{g}$ is generically subject to stringent cosmological limits from the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) [5].
@@
Earlier work has considered clusters of cosmologically stable PBHs in the later universe after matter-radiation equality [20, 21, 22, 23], or focused on PBH mergers in the early Universe without cluster formation (the “quasi-linear regime”) [10, 21, 24, 25, 26, 27, 28, 7, 8, 9], while Refs. [28, 29] explored the formation of heavier PBHs from strongly clustered lighter ones (clusteringenesis), but did not specify a mechanism for this enhanced initial clustering at PBH formation. Ref. [30] studied PBH formation from the growth of density perturbations in PBH domination but did not consider mergers in PBH clusters. Black hole mergers in clusters have also been studied in the context of building supermassive black holes in galaxies [31, 32, 33]. Our analysis is complementary as we consider evaporating PBHs with $m<10^{9}\,\mathrm{g}$ and calculate their clustering behavior in the early universe in a BHD era, which has not been studied before.
@@ III.1 PBH Cluster Formation
$\mu(M)=\begin{cases}\displaystyle 1-\frac{4}{7}\frac{M}{M_{H,i}}&M<M_{H,i}\vskip 3.00003pt\\ \displaystyle\frac{3}{7}{\left(\frac{M}{M_{H,i}}\right)}^{-4/3}&M>M_{H,i}\end{cases}$ (21)
@@ III.2 PBH Binary Capture
where $t_{\mathrm{cap,ins}}$ are respectively the capture and inspiral timescales and the 2b/3b labels refer to two and three body processes. Thus, on average, 2-body interactions lead to a quicker capture, inspiral, and merger than 3-body interactions. As a concrete example, we show in Fig. 2 the rates for 2-body and 3-body capture and inspiral for one choice of free parameters. For the remainder of this work we neglect 3-body interactions, which will not contribute significantly to the PBH merger rate during BHD, and instead focus solely on the 2-body interaction channel.
Figure 2: A comparison of the 2-body capture timescale from Eq. (32) and the average 2-body inspiral timescale from Eq. (35) with the 3-body capture and inspiral timescales from Eqs. (6) and (12) of Ref. [22] for one choice of our free parameters $m_{i}$ and $t_{i}$, showing that Eq. (29) is satisfied for all $N_{i}$.
@@ IV.3 Combined Treatment
$\frac{dM}{dN}=\frac{M}{N}\,\frac{\Gamma_{\mathrm{ev}}+\frac{\Delta m}{{\mkern 2.5mu\overline{\mkern-2.5mum\mkern-2.0mu}\mkern 2.0mu}}\,\Gamma_{\mathrm{mrg}}}{\Gamma_{\mathrm{ev}}+\Gamma_{\mathrm{mrg}}},$ (50)
@@ IV.4 Merged Relics
$N_{\mathrm{col}}^{\mathrm{max}}\approx 10^{10}\left\{\begin{array}[]{l}\displaystyle 12\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!12/7}{\left(\frac{t_{i}}{10^{-20}\,\mathrm{s}}\right)}^{\!-4/7}\vskip 3.00003pt\\ \displaystyle 6.5\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!8/11}\end{array}\right$.\!\!\!, (55)
@@
$m_{\mathrm{relic}}^{\mathrm{max}}\approx 10^{15}\,\mathrm{g}\left\{\begin{array}[]{l}\displaystyle 24\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!3.53}{\left(\frac{t_{i}}{10^{-20}\,\mathrm{s}}\right)}^{\!-1.27}\vskip 3.00003pt\\ \displaystyle 1.4\,{\left(\frac{m_{i}}{10^{9}\,\mathrm{g}}\right)}^{\!1.35}\end{array}\right$.\!\!\!, (56)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\frac{dn_{\mathrm{BH}}}{dm}(\tau)$ $=\int_{1}^{N_{\mathrm{col}}^{\mathrm{max}}}\!dN_{i}\,\frac{dn_{\mathrm{cl}}}{dN_{i}}(\tau)\,\delta\big(m-m_{\mathrm{relic}}(N_{i})\big)$
$\quad+\int_{N_{\mathrm{col}}^{\mathrm{max}}}^{\infty}\!dN_{i}\,N_{i}\frac{dn_{\mathrm{cl}}}{dN_{i}}(\tau)\,\delta\left(m-m_{i}\right)$
@@ IV.5 Important Caveats
  6. 6.
    Non-uniform cluster profile: In Eqs. (37) and (39) we assumed a uniform spherical density profile for PBHs in a cluster. Considering a realistic profile might lead to enhancement of merger rates in the denser core.
  7. 7.
    Primordial binaries: Some binaries may form in the background before clusters can form, by virtue of their initial configuration. These have previously been explored extensively, but Ref. [23] finds some possible implications for gravothermal cluster evolution in a late universe scenario.
  8. 8.
    Monochromatic PBH mass function: Throughout our analysis, we assume that the initial PBH population at $t=t_{i}$ is monochromatic. If PBHs form due to the collapse of primordial density fluctuations, this population corresponds to a large, highly-localized enhancement to the primordial power spectrum. In realistic inflationary models, such features can arise from multi-field dynamics or from temporary departures away from slow-roll evolution (see Ref. [76] for a review). In general, these features can yield a wide range of PBH mass functions, though the details are highly model-dependent. However, if the departure from slow-roll inflation is short compared to the Hubble time during this epoch, then the power spectrum is only modified on a narrow range of scales and the PBH mass function is narrow, as in the scenario we consider.
@@ V.1 Merged Relics as Dark Matter Components
The BHD dynamics outlined in Sec. III predict that for many choices of initial parameters ($m_{i},t_{i}$), there is an appreciable population of merged relics with significantly larger masses than the original PBH population that drives BHD, $m_{\mathrm{relic}}\gg m_{i}$. Consequently, these objects are longer lived and generically survive past BBN to affect the later universe.
There are two qualitatively distinct scenarios for relic masses. For intermediate masses $10^{9}\,\mathrm{g}<m_{\mathrm{relic}}<10^{15}\,\mathrm{g}$, evaporating relics evaporate after BHD and are not constrained by late-time observations. However, because of the energy they inject while evaporating, they do face stringent limits from the successful predictions of standard BBN and CMB cosmology [4, 3]. For $m_{\mathrm{relic}}>10^{15}\,\mathrm{g}$, metastable relics can be cosmologically metastable and account for an appreciable fraction of the dark matter density. This population is constrained by late-time observations including lensing, gravitational waves, and particle fluxes from Hawking evaporation, although there is an “asteroid mass” window between $10^{17}-10^{23}\,\mathrm{g}$ with no current constraints [4, 3].
@@ V.2 Gravitational Waves
where $g_{*,S}$ is the number of entropic degrees of freedom, an RH label denotes a quantity at reheating ($t=\tau$), a 0 label represents a present day quantity, and $m$ is the characteristic mass of the PBHs that undergo runaway mergers. Although the peak frequency in Eq. (61) is far outside of any known observational technique, as noted in Ref. [10], the infrared tail of this stochastic background might be within the reach of the Big Bang Observer [83, 84].
@@ VI Conclusion
However, if the initial PBHs have a shot noise power spectrum, we find that relics cannot accommodate the full dark matter abundance. We find that, by assembling relic PBHs in the viable dark matter window ($m_{\mathrm{relic}}\sim 10^{17}-10^{23}\,\mathrm{g}$) with the observed DM abundance, there is an irreducible tail of evaporating relics at lower masses ($10^{9}-10^{15}\,\mathrm{g}$) whose abundance is excluded by the BBN and CMB limits. Thus, under our minimal assumptions, this mechanism cannot account for any viable present day PBH abundance. While a different ansatz for the initial PBH power spectrum (e.g. more initial clustering than shot noise predicts) might result in enhanced merging, it remains to be seen whether that could allow for a viable dark matter formation scenario. We leave such analyses for future work.
Finally, we note that our results rely on several assumptions about PBH cluster evolution discussed in Sec. IV.5. Notably, we treat all of the PBHs in a cluster as evolving collectively and simultaneously, and we assume that gravitational bremsstrahlung, cluster-cluster interactions, and relativistic effects (among others) do not significantly affect our results. While we have presented plausible arguments for why our results should not be affected in most of these cases, many of these effects are not analytically tractable and may, nonetheless, allow for different conclusions to hold; especially when multiple effects are combined. Thus, in order to robustly characterize cluster evolution and relic formation during BHD, a full $N$-body simulation will ultimately be necessary. Running such a simulation is beyond the scope of the present work, but we strongly encourage future efforts to better understand this simple, yet physically rich cosmology.
@@ Acknowledgments
Data Availability
The code and data to replicate the results of this article are available at [96].
@@ Appendix A Shot Noise Power Spectrum
[…]